OpenStudy (cometailcane):

What is the slope of the line that passes through the given points? (7, -9) and (7, -1)

OpenStudy (cometailcane):

i know the answer is undefined but i don't know how you figure that out

OpenStudy (alexandervonhumboldt2):

(-9+1)/(7-7)=-8/0=undefined slope

OpenStudy (cometailcane):

oh ok